OpenStudy (anonymous):

What is the value of y in the solution to the following system of equations? 5x − 3y = −11 2x − 6y = −14

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

is it 2?

jimthompson5910 (jim_thompson5910):

very good the solution to the system as an ordered pair is (-1,2) focusing on just y we get y = 2

OpenStudy (anonymous):

thanks! Use the substitution method to solve the following system of equations: 3x + 5y = 3 x + 2y = 0 i got (6,-3)?

jimthompson5910 (jim_thompson5910):

You can plug the point back into each equation to see if you get a true result 3x + 5y = 3 3(6) + 5(-3) = 3 ... plug in x = 6 and y = -3 18 - 15 = 3 3 = 3 I'll let you check the other one

OpenStudy (anonymous):

6+2(-3)=0 6+-6=0 0=0

jimthompson5910 (jim_thompson5910):

so (6,-3) is confirmed to be the solution

OpenStudy (anonymous):

this ones really confusing me.. In the below system, solve for y in the first equation. x + 3y = 6 2x − y = 10 i got y=-1/3x+2

jimthompson5910 (jim_thompson5910):

I'm getting \[\Large y = -\frac{1}{3}x + 2\] as well

OpenStudy (anonymous):

awesome!! Miranda bought some apples and oranges at the store for fruit salad. Owen bought three times as many apples and twice as many oranges. Miranda has 5 fruits in her bag, but Owen has 12 fruits in his bag. How many apples did Miranda purchase? I got 2?

jimthompson5910 (jim_thompson5910):

you're on a roll

jimthompson5910 (jim_thompson5910):

you are correct

OpenStudy (anonymous):

Yess!!! last one: Mark, Jessica, and Nate each downloaded music from the same website. Mark downloaded 10 songs in total consisting of pop, rock, and hip hop. Jessica downloaded five times as many pop songs, twice as many rock, and three times as many hip hop songs as Mark. She downloaded 28 songs total. If Nate downloaded 20 songs total with three times as many pop songs, three times as many rock songs, and the same number of hip hop songs as Mark, how many songs of each type did Mark download? I got 1 pop songs, 4 rock and 5 hip hop

jimthompson5910 (jim_thompson5910):

I'm getting the same as well

OpenStudy (anonymous):

omg yes! thank you so much!

jimthompson5910 (jim_thompson5910):

you're welcome